If α, β are the roots of ax 2 +bx+c=0 then, the value of α 2 /β 2 + β 2 /α 2 is



If  α, β are the roots of ax2+bx+c=0 then, the value of α22 + β22 is


Grade:Upto college level

2 Answers

jitender lakhanpal
10 years ago

 α+β= -b/a ------1  
 α β= c/a -------2    given


 (α^4+β^4)/( α β)^4       sum of roots
(α^4+β^4) = {(α+β)^2-2α β }^2-2 (α β)^2} -----by algebraic identities

now put the equations 1 and 2 in above expression

product of roots = 1

we know the sum and product we can easily form the equation

 

We are  here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

 

Thanks and Regards

Jitender

(askIITians expert)

vijay
10 years ago